OpenStudy (anonymous):

solve: 2 absolute value of x minus 3 =7

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

\[2|x-3|=7\]

OpenStudy (mertsj):

Divide both sides by 2 |x-3|=7/2 x-3=7/2 or x-3=-7/2 Solve both equations.

OpenStudy (anonymous):

how do u solve it thats what i dont understand

OpenStudy (mertsj):

Add 3 to both sides of the equations I have written. Can you add 3 to 7/2? Can you add 3 to -7/2?

OpenStudy (anonymous):

no u get a fraction

OpenStudy (mertsj):

Yes. Sometimes in mathematics, you will have answers that are fractions.

OpenStudy (anonymous):

i know but the answers i have are not answers on my worksheet

OpenStudy (mertsj):

3 + 7/2=3+3 1/2=6 1/2=6.5 3-7/2=3-3 1/2=-1/2=-.5

OpenStudy (anonymous):

the answers on my sheet are: x=-10,x=10. x=-2,x=2. x--5,x=5. or x=-8,x=8.

OpenStudy (mertsj):

So if that is the case, I have not interpreted your problem correctly and the equation editor is not working. Is your problem 2 times the absolute value of (x-3)=7 or is it something else?

OpenStudy (anonymous):

my problem is 2 absolute value of x. (x by itself.) minus 3 =7

OpenStudy (mertsj):

Ok. Add 3 to both sides

OpenStudy (mertsj):

Then divide both sides by 2.

OpenStudy (mertsj):

I'm going to restart my browser. Maybe things will work better.

OpenStudy (mertsj):

\[2|x|-3=7\]

OpenStudy (mertsj):

\[|x|=5\]

OpenStudy (anonymous):

yes that is what it looks like

OpenStudy (mertsj):

x=5 or x = -5

OpenStudy (anonymous):

thank u im sorry for the confusion :(

OpenStudy (mertsj):

yw. It's so much easier when the equation editor is working.

OpenStudy (anonymous):

ikr

OpenStudy (mertsj):

What is ikr?

OpenStudy (anonymous):

how do u get the equation editor

OpenStudy (anonymous):

i know right

OpenStudy (mertsj):

That blue button below this box that says "equation".

OpenStudy (anonymous):

oh.. ok ... thanks

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

will come back at school have to go to library. hang on

OpenStudy (mertsj):

ok